# Generalized Lanczos method for systematic optimization of neural-network quantum states

### v1

@@ Abstract
Recently, artificial intelligence for science has made significant inroads into various fields of natural science research. In the field of quantum many-body computation, researchers have developed numerous ground state solvers based on neural-network quantum states (NQSs), achieving ground state energies with accuracy comparable to or surpassing traditional methods such as variational Monte Carlo methods, density matrix renormalization group, and quantum Monte Carlo methods. Here, we combine supervised learning, reinforcement learning, and the Lanczos method to develop a systematic approach to improving the NQSs of many-body systems, which we refer to as the NQS Lanczos method. The algorithm mainly consists of two parts: the supervised learning part and the reinforcement learning part. Through supervised learning, the Lanczos states are represented by the NQSs. Through reinforcement learning, the NQSs are further optimized. We analyze the reasons for the underfitting problem and demonstrate how the NQS Lanczos method systematically improves the energy in the highly frustrated regime of the two-dimensional Heisenberg $J_{1}$-$J_{2}$ model. Compared to the existing method that combines the Lanczos method with the restricted Boltzmann machine, the primary advantage of the NQS Lanczos method is its linearly increasing computational cost.
neural-network quantum states, the Lanczos method, supervised learning, reinforcement learning
@@ I Introduction
In this paper, we propose an alternative implementation of the Lanczos method, called the NQS Lanczos method, which consists of two parts: the supervised-learning Lanczos (SLL) and the reinforcement-learning Lanczos (RLL) algorithms. The SLL algorithm further consists of two parts: supervised training and diagonalization. Through supervised training, the SLL algorithm represents Lanczos states with NQSs, thereby avoiding the calculation of the expectations $\langle H^{2i+1}\rangle$. When the supervised learning is terminated, a set of basis NQSs are constructed. Then a Hamiltonian matrix is constructed with these basis NQSs. Diagonalizing the Hamiltonian matrix, a superposition state is obtained. The RLL algorithm uses a reinforcement learning scheme for the amplitude part of the superposition state. When the RLL algorithm is terminated, an optimized superposition state and the improved energy are obtained. Using the optimized superposition state as a new state for the next iteration, the full loop of the NQS Lanczos method is achieved by repeating the procedure.
@@
The organization of this paper is as follows. In Section II, we first introduce the structure of the NQSs used for supervised learning. Next, we show the details of the supervised learning, including the loss functions and the optimization strategy employed. Following this, we introduce a reinforcement learning scheme specifically designed for the amplitude component. Finally, building on the aforementioned steps, we present the complete procedure of the NQS Lanczos method. In Section III, we show the calculated results for the two-dimensional Heisenberg $J_{1}$-$J_{2}$ model and provide an analysis and discussion. Lastly, in Section IV, we discuss the strengths and limitations of the NQS Lanczos method and provide an outlook for future improvements.
@@ II Method
In this section, we introduce a representation of the NQS. Based on this representation, we introduce the details of the SLL algorithm and design the loss function and optimization strategy. To address issues arising from underfitting in the SLL algorithm, we propose an RLL scheme to optimize the states further. The section concludes with an overview of the complete NQS Lanczos method.
@@ II.1 Neural-network quantum states
II.2 The supervised-learning Lanczos algorithm
The Lanczos method [31] was first introduced by Cornelius Lanczos in 1952. In 2013, Hu et al. [3] used this approach to further improve the states obtained by VMC in the calculation of many-body systems, achieving promising results at that time.
In this work, we use the Lanczos state (denoted as $|\psi_{i}\rangle$) at Lanczos step $i$ as the target of supervised learning. Through supervised learning, the sign network and amplitude network of the NQS are optimized so that the NQS can be used as an approximate representation of $|\psi_{i}\rangle$. The NQS and supervised learning are combined to implement the Lanczos procedure. The procedure is as follows.
i. Give an initial state $|\psi_{0}\rangle$. This state can be arbitrary (which is not orthogonal to the ground state). In this work, we use aCNN as the initial state, denoted as $|\psi^{\text{net}}_{0}\rangle$.
ii. Define $|v_{1}\rangle=H|\psi^{\text{net}}_{0}\rangle-a_{0}|\psi^{\text{net}}_{0}\rangle$, where $a_{0}={\langle\psi^{\text{net}}_{0}|H|\psi^{\text{net}}_{0}\rangle}/{\langle\psi^{\text{net}}_{0}|\psi^{\text{net}}_{0}\rangle}$ is the expectation of the Hamiltonian on the state $|\psi^{\text{net}}_{0}\rangle$. Let $b_{1}|\psi^{\mathrm{trg}}_{1}\rangle=|v_{1}\rangle$, where $b_{1}=\sqrt{{\langle v_{1}|v_{1}\rangle}/{\langle\psi^{\text{net}}_{0}|\psi^{\text{net}}_{0}\rangle}}$ is the normalization factor that normalizes $|v_{1}\rangle$ with respect to $|\psi^{\text{net}}_{0}\rangle$, namely $\langle\psi^{\mathrm{trg}}_{1}|\psi^{\mathrm{trg}}_{1}\rangle=\langle\psi^{\text{net}}_{0}|\psi^{\text{net}}_{0}\rangle$. The Lanczos state $|\psi^{\mathrm{trg}}_{1}\rangle$ is set as the target state to be learned. Through Monte Carlo sampling, the coefficients $a_{0}$ and $b_{1}$ can be calculated. Then the target state $|\psi^{\mathrm{trg}}_{1}\rangle$ can be expressed, with the coefficients $a_{0}$, $b_{1}$, and the initial state $|\psi_{0}\rangle$.
iii. Take a configuration $\sigma$ as a training sample for supervised learning. The sign and modulus of $\psi^{\mathrm{trg}}_{1}(\sigma)$ are taken as labels for the supervised learning of the sign network and the amplitude network, respectively. Then, the sign network $\text{SNet}_{1}$ and the amplitude network $\text{ANet}_{1}$ can be trained through supervised learning. Once the losses converge, the wave function $\psi^{\text{net}}_{1}=S_{1}\cdot e^{A_{1}}$ of the Lanczos state is obtained, that is, the Lanczos state is represented by a NQS.
iv. Define
$|v_{i+1}\rangle=H|\psi^{\text{net}}_{i}\rangle-a_{i}|\psi^{\text{net}}_{i}\rangle-b_{i}|\psi^{\text{net}}_{i-1}\rangle,$ (1)
@@ II.2 The supervised-learning Lanczos algorithm
$a_{i}$ $=\frac{\langle\psi^{\text{net}}_{i}|H|\psi^{\text{net}}_{i}\rangle}{\langle\psi^{\text{net}}_{i}|\psi^{\text{net}}_{i}\rangle}.$ (2)
The Lanczos state $|\psi_{i+1}^{\mathrm{trg}}\rangle$ is constructed as
$|\psi_{i+1}^{\mathrm{trg}}\rangle$ $=\frac{|v_{i+1}\rangle}{b_{i+1}},$ (3)
@@
$b_{i+1}=\sqrt{\frac{\langle v_{i+1}|v_{i+1}\rangle}{\langle\psi^{\text{net}}_{0}|\psi^{\text{net}}_{0}\rangle}}.$ (4)
The coefficients $a_{i}$ and $b_{i+1}$ are calculated through Monte Carlo sampling.
v. Train a sign network $\text{SNet}_{i}$ and an amplitude network $\text{ANet}_{i}$ through supervised learning. The wave function $\psi^{\text{net}}_{i+1}=S_{i+1}\cdot e^{A_{i+1}}$ at Lanczos step $i$ is obtained.
@@
can be obtained, where $c_{i}$ is the superposition coefficient, and $p$ is the number of Lanczos steps. Details of the diagonalization process are given in Appendix A.
II.3 Loss function and optimization
@@
The goal of supervised learning of the amplitude network is to make the distribution described by $|\psi^{\mathrm{net}}_{i}|^{2}$ consistent with the distribution described by $|\psi^{\mathrm{trg}}_{i}|^{2}$. In deep learning field, the Kullback-Leibler (KL) divergence is often used to measure the difference between two distributions. However, in our tests, the KL divergence loss function performed poorly. We choose to use the loss function based on the mean squared error, $\mathbb{E}_{x\sim p}||y_{x}-f(x)||$. It is to minimize the difference squared $\left(|\psi^{\text{trg}}_{i}|-|\psi^{\text{net}}_{i}|\right)^{2}$ between the target and the prediction. To achieve this, samples should be obtained from the distributions described by both $|\psi^{\text{trg}}_{i}|^{2}$ and $|\psi^{\text{net}}_{i}|^{2}$. The loss function of the amplitude part is
@@
It is worth noting that the probability distribution $p^{\text{trg}}_{i}(\sigma)$ and the label $\psi^{\text{trg}}_{i}(\sigma)$ are both determined by the previous NQSs, whose parameters are fixed and are not involved in optimization here. While the probability distribution $p^{\text{net}}_{i}(\sigma,\theta)$ contains the parameters $\theta$ to be optimized. The gradient with respect to these parameters must be considered. According to the theory of automatic differentiation Monte Carlo (ADVMC) [43], the expectation of an observable $O(\sigma,\theta)$ is given by
@@
where $\perp(\theta)$ is a function defined as $\perp(\theta)=\theta$ in forward propagation, and ${\partial\perp(\theta)}/{\partial\theta}=0$ in backward propagation. This reformulation applies to both normalized and unnormalized probability distributions. The loss function for the amplitude network can be rewritten as
@@
II.4 The reinforcement-learning Lanczos algorithm
In the tests of aCNN [35], the energies achieved by optimizing only the amplitude part of the wave function are more accurate than those of the complex-valued CNN [44] with the same number of parameters. This inspires us to further optimize the amplitude of the superposition of NQSs to mitigate the negative effects of insufficient convergence in the supervised learning. We select an amplitude network $\text{ANet}_{j}$ with $j\neq 0$ to do a reinforcement learning, while fixing all the parameters of the other networks as well as the superposition coefficients.
During the reinforcement learning, we use energy as the loss function and employ the ADVMC method [43] to calculate the derivative of energy with respect to the network parameters. The popular Adam algorithm [45] is used to update the parameters. The energy is expressed as
$E$ =$H\right>=\langle E_{\mathrm{loc}}(\sigma,\theta_{i})\rangle_{p(\sigma,\theta_{i})}$
@@ II.4 The reinforcement-learning Lanczos algorithm
[…]
The reinforcement learning effectively optimizes the amplitude network $\text{ANet}_{j}$ further. The goal is not to make $|\psi^{\mathrm{net}}_{j}|$ match $|\psi^{\mathrm{trg}}_{j}|$ but to make the overall superposition state $|\Psi\rangle$ approach the true ground state. After optimizing each $\text{ANet}_{j}$ with $0<j\leq p$ (note that $\text{ANet}_{0}$ has already been obtained through amplitude optimization and does not need further optimization), we obtain a new state $|\tilde{\Psi}\rangle$. This new state can then be used as the updated $|\psi^{\mathrm{net}}_{0}\rangle$ for the next NQS Lanczos loop to further improve the results.
II.5 The NQS Lanczos method
Figure 3: Flowchart of the NQS Lanczos method. The method begins with $|\psi^{\mathrm{net}}_{0}\rangle$ as a starting point and ultimately outputs the improved energy $\tilde{E}$ and state $|\tilde{\Psi}\rangle$. $|\tilde{\Psi}\rangle$ can be used as a new $|\psi^{\mathrm{net}}_{0}\rangle$ for the next NQS Lanczos loop. The NQS Lanczos method consists of two parts, the SLL algorithm and the RLL algorithm. The SLL algorithm further consists of two parts: supervised learning of Lanczos states and diagonalization of the Hamiltonian matrix. The output of the SLL algorithm consists of the superposition state $|\Psi\rangle$ and the improved energy $E$. The RLL algorithm further optimizes $|\Psi\rangle$ and gives the final output of the NQS Lanczos method. The number of the Lanczos steps is denoted as $p$. The matrices $M$ and $H$ are used to produce the superposition state, as shown in Appendix A.
In this section, we outline the entire procedure of the NQS Lanczos method, as shown in Fig. 3. The NQS Lanczos method consists of two parts, the SLL algorithm and the RLL algorithm. The first step is to provide an initial state $|\psi^{\mathrm{net}}_{0}\rangle$, based on which the coefficients $a_{i}$ and $b_{i+1}$ are calculated. Then the target $|\psi^{\mathrm{trg}}_{i}\rangle$ is constructed, according to which, samples and labels are generated for supervised learning. Through the supervised learning, both the sign and the amplitude networks are optimized. The wave function $\psi^{\mathrm{net}}_{i}=S_{i}e^{A_{i}}$ represented by NQS is obtained. As the Lanczos step $i$ goes, a set $\{|\psi^{\mathrm{net}}_{i}\rangle\}$ ($i=0,\dots,p$) is obtained.

### v2

@@ Abstract
Recently, artificial intelligence for science has made significant inroads into various fields of natural science research. In the field of quantum many-body computation, researchers have developed numerous ground state solvers based on neural-network quantum states (NQSs), achieving ground state energies with accuracy comparable to or surpassing traditional methods such as variational Monte Carlo methods, density matrix renormalization group, and quantum Monte Carlo methods. Here, we combine supervised learning, variational Monte Carlo (VMC), and the Lanczos method to develop a systematic approach to improving the NQSs of many-body systems, which we refer to as the NQS Lanczos method. The algorithm mainly consists of two parts: the supervised learning part and the VMC optimization part. Through supervised learning, the Lanczos states are represented by the NQSs. Through VMC, the NQSs are further optimized. We analyze the reasons for the underfitting problem and demonstrate how the NQS Lanczos method systematically improves the energy in the highly frustrated regime of the two-dimensional Heisenberg $J_{1}$-$J_{2}$ model. Compared to the existing method that combines the Lanczos method with the restricted Boltzmann machine, the primary advantage of the NQS Lanczos method is its linearly increasing computational cost.
neural-network quantum states, the Lanczos method, supervised learning, VMC
@@ I Introduction
In this paper, we propose an alternative implementation of the Lanczos method, called the NQS Lanczos method, which consists of two parts: the supervised-learning Lanczos (SLL) algorithm and the VMC optimization part. The SLL algorithm further consists of two parts: supervised training and diagonalization. Through supervised training, the SLL algorithm represents Lanczos states with NQSs, thereby avoiding the calculation of the expectations $\langle H^{2i+1}\rangle$. When the supervised learning is terminated, a set of basis NQSs are constructed. Then a Hamiltonian matrix is constructed with these basis NQSs. Diagonalizing the Hamiltonian matrix, a superposition state is obtained. The VMC optimization part are used to optimize the amplitude part of the superposition state. When the VMC optimization part is terminated, an optimized superposition state and the improved energy are obtained. Using the optimized superposition state as a new state for the next iteration, the full loop of the NQS Lanczos method is achieved by repeating the procedure.
@@
The organization of this paper is as follows. In Section II, we first introduce the structure of the NQSs used for supervised learning. Next, we show the details of the supervised learning, including the loss functions and the optimization strategy employed. Following this, we introduce a VMC optimization scheme specifically designed for the amplitude component. Finally, building on the aforementioned steps, we present the complete procedure of the NQS Lanczos method. In Section III, we show the calculated results for the two-dimensional Heisenberg $J_{1}$-$J_{2}$ model and provide an analysis and discussion. Lastly, in Section IV, we discuss the strengths and limitations of the NQS Lanczos method and provide an outlook for future improvements.
@@ II Method
In this section, we introduce a representation of the NQS. Based on this representation, we introduce the details of the SLL algorithm and design the loss function and optimization strategy. To address issues arising from underfitting in the SLL algorithm, we introduce a VMC optimization scheme to optimize the states further. The section concludes with an overview of the complete NQS Lanczos method.
@@ II.1 Neural-network quantum states
II.2 The Lanczos algorithm
The Lanczos method [31] was first introduced by Cornelius Lanczos in 1952. In 2013, Hu et al. [3] used this approach to further improve the states obtained by the VMC method in the simulation of many-body systems, achieving promising results at that time. The procedure of the Lanczos algorithm is as follows.
i. Give an initial state $|\psi_{0}\rangle$. This state, which is not orthogonal to the ground state, can be arbitrary.
ii. Define $|v_{1}\rangle=H|\psi_{0}\rangle-a_{0}|\psi_{0}\rangle$, where $a_{0}={\langle\psi_{0}|H|\psi_{0}\rangle}/{\langle\psi_{0}|\psi_{0}\rangle}$ is the expectation of the Hamiltonian on the state $|\psi_{0}\rangle$. Let $b_{1}|\psi_{1}\rangle=|v_{1}\rangle$, where $b_{1}=\sqrt{{\langle v_{1}|v_{1}\rangle}/{\langle\psi_{0}|\psi_{0}\rangle}}$ is the normalization factor that normalizes $|v_{1}\rangle$ with respect to $|\psi_{0}\rangle$, namely $\langle\psi_{1}|\psi_{1}\rangle=\langle\psi_{0}|\psi_{0}\rangle$.
iii. Define
$|v_{i+1}\rangle=H|\psi_{i}\rangle-a_{i}|\psi_{i}\rangle-b_{i}|\psi_{i-1}\rangle,$ (1)
@@ II.2 The Lanczos algorithm
$a_{i}$ $=\frac{\langle\psi_{i}|H|\psi_{i}\rangle}{\langle\psi_{i}|\psi_{i}\rangle}.$ (2)
The Lanczos state $|\psi_{i+1}\rangle$ is constructed as
$|\psi_{i+1}\rangle$ $=\frac{|v_{i+1}\rangle}{b_{i+1}},$ (3)
@@
$b_{i+1}=\sqrt{\frac{\langle v_{i+1}|v_{i+1}\rangle}{\langle\psi_{0}|\psi_{0}\rangle}}.$ (4)
Step iii can be used for all subsequent iterations.
iv. After $p$ iterations, a set of $p+1$ Lanczos vectors $|\psi_{i}\rangle$, $i=0,\cdots,p$, can be obtained, which forms an orthonormal basis set spanning a reduced Hilbert space. With this basis the Hamiltonian is represented as a tridiagonal matrix,
$T=\begin{bmatrix}a_{0}&b_{1}&0&\cdots&0\\ b_{1}&a_{1}&b_{2}&\cdots&0\\ 0&b_{2}&a_{2}&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\cdots&a_{p}\\ \end{bmatrix}$.
v. Diagonalize $T$ to obtain eigenvalues and corresponding eigenvectors.
II.3 The supervised-learning Lanczos algorithm
In this work, we use the Lanczos state $|\psi_{i}\rangle$ at Lanczos step $i$ as the target of supervised learning. Through supervised learning, the sign network and amplitude network of the NQS are optimized so that the NQS can be used as an approximate representation of $|\psi_{i}\rangle$. The NQS and supervised learning are combined to implement the Lanczos procedure.
First, we use aCNN as the initial state, denoted as $|\psi^{\text{net}}_{0}\rangle$. Then the Lanczos state $|\psi_{i}\rangle$ at step $i$ is set as the target state, denoted as $|\psi^{\mathrm{trg}}_{i}\rangle$, to be learned. Through Monte Carlo sampling, the coefficients $a_{i}$ and $b_{i+1}$ can be calculated. Then the target state $|\psi^{\mathrm{trg}}_{i}\rangle$ can be expressed with these coefficients and the previrous states.
Take a configuration $\sigma$ as a training sample for supervised learning. The sign and amplitude of $\psi^{\mathrm{trg}}_{i}(\sigma)$ are taken as labels for the supervised learning of the sign network and the amplitude network, respectively. Then, the sign network $\text{SNet}_{i}$ and the amplitude network $\text{ANet}_{i}$ can be trained through supervised learning. Once the losses converge, the wave function $\psi^{\text{net}}_{i}=S_{i}\cdot e^{A_{i}}$ of the Lanczos state at step $i$ is obtained, that is, the Lanczos state is represented by a NQS.
@@ II.3 The supervised-learning Lanczos algorithm
can be obtained, where $c_{i}$ is the superposition coefficient. Details of the diagonalization process are given in Appendix A.
II.4 Loss function and optimization
@@
[…]
@@
It is worth noting that the probability distribution $p^{\text{trg}}_{i}(\sigma)$ and the label $\psi^{\text{trg}}_{i}(\sigma)$ are both determined by the previous NQSs, whose parameters are fixed and are not involved in optimization here. While the probability distribution $p^{\text{net}}_{i}(\sigma,\theta)$ contains the parameters $\theta$ to be optimized. The gradient with respect to these parameters must be considered. According to the theory of automatic differentiation Monte Carlo (ADMC) [43], the expectation of an observable $O(\sigma,\theta)$ is given by
@@
where $\perp(x)$ is the detach function, which was first introduced in Section III A of Ref. [43], features $\perp(x)=x$ in forward propagation and ${\partial\perp(x)}/{\partial x}=0$ in backward propagation. $x$ does not propagate gradients in the automatic differentiation process. This reformulation applies to both normalized and unnormalized probability distributions. The loss function for the amplitude network can be rewritten as
@@
The loss defined in Eq. (6) is the MSE loss. Another choice is to adopt the fidelity between the quantum states as the loss function, as employed in Refs. [44, 45, 46]. The loss function based on fidelity is given by
$L(\psi^{\mathrm{net}}_{i}(\sigma,\theta),\psi^{\mathrm{trg}}_{i}(\sigma))$ $=-\log\left[\frac{\left|\langle\psi^{\mathrm{net}}_{i}(\sigma,\theta)|\psi^{\mathrm{trg}}_{i}(\sigma)\rangle\right|}{\left|{|\psi^{\mathrm{net}}_{i}(\sigma,\theta)\rangle}\right|\left|{|\psi^{\mathrm{trg}}_{i}(\sigma)\rangle}\right|}\right]$
$=-\log\sqrt{\alpha\beta},$ (11)
where
$\alpha$ $=\left|\left<\frac{\psi^{\mathrm{trg}}_{i}(\sigma)}{\psi^{\mathrm{net}}_{i}(\sigma,\theta)(\sigma)}\right>_{\psi^{\mathrm{net}}_{i}(\sigma,\theta)}\right|,$
$\beta$ $=\left|\left<\frac{\psi^{\mathrm{net}*}_{i}(\sigma,\theta)(\sigma)}{\psi^{\mathrm{trg}*}_{i}(\sigma)}\right>_{\psi^{\mathrm{trg}}_{i}(\sigma)}\right|.$ (12)
The derivative of the loss function based on fidelity with respect to the parameter $\theta$ can be referred to Ref. [46]. However, the optimization of this loss function is difficult due to its inherent numerical instability, which often leads to overflow during training. To address this, we have implemented two distinct strategies. Strategy 1 (denoted ad s1) introduces a regularization term to the loss function, constraining the amplitudes within an appropriate range. Strategy 2 (denoted as s2) involves dynamic monitoring during optimization. If the components $\alpha$ or $\beta$ in the loss function exhibit anomalous behavior, the corresponding term is detached during backpropagation and is temporarily excluded from optimization until its values return to a stable, reasonable range.
While these strategies involve explicit manual design, they often serve as effective auxiliary treatments that stabilize the optimization process in our tests. A comparative analysis of the convergence behavior for different loss functions is provided in Section III.B. In subsequent numerical experiments, the MSE loss demonstrated robust and competitive performance. Consequently, unless stated otherwise, we choose the MSE loss as the default throughout the rest of this paper.
@@
II.5 VMC on the superposition state
In the tests of aCNN [35], the energies achieved by optimizing only the amplitude part of the wave function are more accurate than those of the complex-valued CNN [47] with the same number of parameters. This inspires us to further optimize the amplitude of the superposition of NQSs to mitigate the negative effects of insufficient convergence in the supervised learning. We select an amplitude network $\text{ANet}_{j}$ with $j\neq 0$ to do a VMC optimization, while fixing all the parameters of the other networks as well as the superposition coefficients.
During the VMC optimization, we use energy as the loss function and employ the ADMC method [43] to calculate the derivative of energy with respect to the network parameters. The popular Adam algorithm [48] is used to update the parameters. The energy is expressed as
$E$ $=\left<H\right>=\langle E_{\mathrm{loc}}(\sigma,\theta_{i})\rangle_{p(\sigma,\theta_{i})}$
@@ II.5 VMC on the superposition state
is the probability distribution, and $\langle\cdot\rangle$ represents the expectation.
The VMC optimization part effectively optimizes the amplitude network $\text{ANet}_{j}$ further. The goal is not to make $|\psi^{\mathrm{net}}_{j}|$ match $|\psi^{\mathrm{trg}}_{j}|$ but to make the overall superposition state $|\Psi\rangle$ approach the true ground state. After optimizing each $\text{ANet}_{j}$ with $0<j\leq p$ (note that $\text{ANet}_{0}$ has already been obtained through amplitude optimization and does not need further optimization), we obtain a new state $|\tilde{\Psi}\rangle$. This new state can then be used as the updated $|\psi^{\mathrm{net}}_{0}\rangle$ for the next NQS Lanczos loop to further improve the results.
II.6 The NQS Lanczos method
Figure 3: Flowchart of the NQS Lanczos method. The method begins with $|\psi^{\mathrm{net}}_{0}\rangle$ as a starting point and ultimately outputs the improved energy $\tilde{E}$ and state $|\tilde{\Psi}\rangle$. $|\tilde{\Psi}\rangle$ can be used as a new $|\psi^{\mathrm{net}}_{0}\rangle$ for the next NQS Lanczos loop. The NQS Lanczos method consists of two parts, the SLL algorithm and the VMC optimization part. The SLL algorithm further consists of two parts: supervised learning of Lanczos states and diagonalization of the Hamiltonian matrix. The output of the SLL algorithm consists of the superposition state $|\Psi\rangle$ and the improved energy $E$. The VMC optimization part further optimizes $|\Psi\rangle$ and gives the final output of the NQS Lanczos method. The number of the Lanczos steps is denoted as $p$. The matrices $M$ and $H$ are used to produce the superposition state, as shown in Appendix A.
In this section, we outline the entire procedure of the NQS Lanczos method, as shown in Fig. 3. The NQS Lanczos method consists of two parts, the SLL algorithm and the VMC optimization part. The first step is to provide an initial state $|\psi^{\mathrm{net}}_{0}\rangle$, based on which the coefficients $a_{i}$ and $b_{i+1}$ are calculated. Then the target $|\psi^{\mathrm{trg}}_{i}\rangle$ is constructed, according to which, samples and labels are generated for supervised learning. Through the supervised learning, both the sign and the amplitude networks are optimized. The wave function $\psi^{\mathrm{net}}_{i}=S_{i}e^{A_{i}}$ represented by NQS is obtained. As the Lanczos step $i$ goes, a set $\{|\psi^{\mathrm{net}}_{i}\rangle\}$ ($i=0,\dots,p$) is obtained.